OpenStudy (anonymous):

Solve using the substitution method. 1/2x-1/5y = 7, 1/6x+ 2/3= -4

OpenStudy (anonymous):

take 1/x = u and 1/y = v, then equations become u/2-v/5=7 u/6+2/3=-4.. solve these with normal method then set 1/x = u and solve for x and y :D

hero (hero):

Solve both equations for y.. Set y = y and substitute the equations in terms of x in their place, then solve for x

OpenStudy (anonymous):

Is this right? 1/2x-1/5y = 7, ---1) 1/6x+ 2/3= -4 ---2) Rearrange equation 2) 1/6x+ 2/3= -4 1/6x = -4 -2/3 x = -28 substitute x = -28 into 1) 1/2 x -28 -1/5y =7 -14 - 1/5y = 7 1/5y = -14 -7 y = 105